An initial amount of £6000 is your savings account. The bank is giving you interest at the rate of 10% per year (annum). After n years the amount of money stored in your bank account will be £6000*(1.10)^n, where n is the
number of years. Evaluate after how many years you will
have £22000 deposited in your bank account?

Answer by Boreal(15235)

22000=6000(1.1)^n
11/3=1.1^n, dividing by 6000 both sides and reducing fractions
ln both sides
ln (11/3)=n ln (1.1)
taking the ln (11/3) and dividing by ln (1.1)
n=13.63 years
Check.
6000*(1.1)^13.63=$21,995.49, within rounding error.